OpenStudy (anonymous):

3d+4=2+3d-1/2 --solve and is the equation an identity or no solution.

OpenStudy (anonymous):

well... all you need to do is: 3d + 4 = 3d +1.5 3d-3d +4= 3d-3d+1.5 4 = 1.5 ( false) so your answer should be : No solutions ( 4 > 1.5, not = to 1.5) :)